How long would it take 4.0 x 10²⁰ atoms to decay to 1.0 x 10¹⁹ atoms if their half-life was 14.7 years?

1
Expert's answer
2022-06-06T08:55:16-0400
"N=N_02^{-t\/t_{1\/2}}""1.0*10^{19}=4.0*10^{20}*2^{-t/14.7\:\rm yrs}""t=-14.7*\log_2(1.0*10^{19}/4.0*10^{20})=78\:\rm yrs"